A ball of mass 0.564 kg moving east (+xdirection) with a speed of 3.70 m/s collides head-on with a 0.282 kg ball at rest. Assume that the collision is perfectly elastic.

A) What is be the speed of the 0.564-kg ball after the collision?

Express your answer to three significant figures and include the appropriate units.

B) What is the speed of the 0.282-kg ball after the collision?

Express your answer to three significant figures and include the appropriate units.
Answer #1

here,

mass of ball 1 , m1 = 0.564 kg

mass of ball 2 , m2 = 0.282 kg

initial speed of ball 1 , u1 = 3.7 m/s

let the final speed of balls be v1 and v2

using conservation of momentum

m1 * u1 = m1 * v1 + m2 * v2

0.564 * 3.7 = 0.564 * v1 + 0.282 * v2 ......(1)

and

using conservation of kinetic energy

0.5 * m1 * u1^2 = 0.5 * m1 * v1^2 + 0.5 * m2 * v2^2

0.564 * 3.7^2 = 0.564 * v1^2 + 0.282 * v2^2 ......(2)

from (1) and (2)

v1 = 1.23 m/s

v2 = 4.93 m/s

A)

the speed of the 0.564 kg ball is 1.23 m/s

B)

the speed of 0.282 kg ball is 4.93 m/s